OpenStudy (anonymous):

Determine f(2)(x) when f(x) =sin x/cos x + 1

zepdrix (zepdrix):

That notation doesn't make much sense. Do they want you to evaluate f(2)? Because that would make sense :)

OpenStudy (anonymous):

yes they do

zepdrix (zepdrix):

So you just plug a 2 into every place you see an x in the problem. Hmmm 2 is not a special reference angle. So you can't get an exact answer. Are you allowed to use a calculator for this problem?

OpenStudy (anonymous):

sorry i typed that wrong they want just the derivative of f(x)=sinx/cosx+ 1

zepdrix (zepdrix):

Oh I see :) Hmmm looks like we have the quotient rule. Remember how to set up the quotient rule?\[(\frac{ u }{ v })'=\frac{ u'v-uv' }{ v^2 }\] So in this problem: u=sinx v=cosx

zepdrix (zepdrix):

Oh maybe I'm misreading it. Is it (cosx+1) on the bottom?

OpenStudy (anonymous):

yes........

zepdrix (zepdrix):

Ok then your v=cosx + 1 :)

OpenStudy (anonymous):

ok...

zepdrix (zepdrix):

Uh ohhhh I'm getting dots :D You confused lil sandy? :D

OpenStudy (anonymous):

what is the answer